A 24-g sample of radioactuve iodine decays in such a way that the mass remaining after t days is given by
m(t)=24e^-0.089t
where m(t) is measured in grams. After how many dYs is there only 17 g remaining?

_______days
Answer #1

m(t) = 24e-0.089t

Now we have to solve for

m(t) = 17

24e-0.089t = 17

e-0.089t = 17/24

Taking natural log both the sides

ln e-0.089t = ln (17/24)

-0.089t ln e = ln (17/24)

-0.089t = ln (17/24) Since, ln e = 1

-0.089t = -0.34484

t = -0.34484 /  -0.089

t = 3.87 Days Answer

Hope This Helps!